# The Orientation Boundary for Step-Duplicating Recursors: Mechanized Impossibility, Escape, and Certification

### 19 Mar 2026 v7

@@ Abstract
We formalize the orientation boundary for first-order step-duplicating recursors. At the schema, no measure in twelve formalized direct classes can strictly orient the duplicating step uniformly. The classes range from additive and transparent-compositional (excluded unconditionally) through affine, restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized degree-bounded polynomial, max-plus, and tracked vector / pair families (excluded under stated pump or growth hypotheses). Two cross-cutting meta-results tighten the stack: a scalar-projection meta-theorem for the matrix-side barriers and a symbolic variable-condition barrier yielding a KBO-style impossibility corollary for the concrete witness system studied later. To our knowledge, this is the first mechanized direct impossibility result for named termination-measure classes on a fixed terminating system. All internal proofs are formalized in Lean 4.
The boundary is explanatory, not only negative. An escape trichotomy for the explicit direct universe, extended further by scalar-projection representability on the matrix side, a transparency-essentiality theorem with a concrete witness, a dependency-pair escape characterization, a generalized polynomial boundary theorem isolating failure of frozen base dominance, computable barrier-witness extractors, ablations, and SCC synchronization theorems classify how successful methods cross the boundary. Every method that orients the duplicating step does so by importing structural assumptions outside the formalized barrier classes.
The concrete witness calculus KO7 carries a two-layer formal development. For the guarded fragment we internally certify strong normalization, confluence, a certified normalizer, decidable reachability to safe normal-form targets, an exact DM-component ordinal calibration at $\omega^{\omega}$, and a triple-lex certificate image below $\omega^{\omega}\cdot 2$. For the full unguarded system we separately prove root termination by a nonlinear polynomial witness and by a KO7-specialized MPO, lift the nonlinear witness to context-closed strong normalization with an explicit derivation-length bound, and connect the Lean development to TTT2/CeTA through a checked TPDB export and verifier bridge.
@@ 1 Introduction
This paper formalizes both sides of this boundary. We prove impossibility theorems at the schema for twelve axiomatized direct measure classes (enumerated below), classify the escape mechanisms, and validate the boundary on a concrete system, KO7, carrying a full certification chain. These twelve classes are not ad hoc: they are the direct scalar and tracked-tuple fragments of the broader tuple-interpretation taxonomy discussed later in Section 4.
@@
The observation that additive and KBO-style direct measures cannot orient duplicating rules is classical: KBO cannot orient any duplicating rule because of the variable condition [3], and the dependency pair framework was introduced to handle systems where simplification orderings fail [1]. Middeldorp and Zantema [18] established the hierarchy between simple termination and full termination. Our contribution is to axiomatize a broader transparency-constrained compositional class together with a separate affine/linear constructor-local class, and then prove barrier theorems, an escape characterization, a transparency-essentiality theorem, and a mechanized ablation for those classes. To our knowledge, no prior work formalizes impossibility theorems for named termination-measure classes on a fixed terminating system: every member of a specified class fails on that system. Mitterwallner, Middeldorp, and Thiemann [19, 27] formalized in Isabelle/HOL the undecidability of orientation by linear polynomial interpretations and by KBO with subterm coefficients. That is a result about the decision problem across all inputs. Their work does not produce standalone impossibility theorems for named systems; ours does. The broader formalization ecosystem (IsaFoR/CeTA, CoLoR) certifies positive termination proofs but not impossibility theorems for measure classes.
The closest neighboring lines of work therefore answer different questions. The LICS/AFP undecidability line asks whether a method class is algorithmically decidable across all inputs; certificate ecosystems such as IsaFoR/CeTA and CoLoR ask whether a successful external termination argument can be reconstructed and checked. Our claim is narrower and different: for one fixed terminating system, we prove object-level impossibility theorems for explicit direct measure families, explain the escape routes, and package constructive witnesses and classifiers around that boundary. This difference is not only mathematical but architectural. The artifact is not another generic certificate checker. It internalizes a negative boundary theory for a fixed terminating system and then reconnects that Lean development back to standard TRS tooling through a checked TPDB export and verifier bridge.
Correspondingly, the paper does not claim that KO7 defeats termination methods in general. The no-go theorems concern twelve direct measure families and an adjacent direct symbolic barrier. Modular methods and fixed-signature path-order arguments remain on the successful side of the boundary, and the artifact proves exactly that through the DP escape, the nonlinear full-step witness $W$, and the KO7-specialized MPO.
@@
A. The orientation barrier (primary contribution). The main theorem package is a twelve-class direct boundary for the duplicating step, with the latter ten classes blocked under their stated pump or growth hypotheses. The classes fall into three blocks: additive and transparency-constrained compositional measures; affine and nonlinear scalar extensions (restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized degree-bounded polynomial, and max-plus); and four tracked vector/pair families (fixed-dimension componentwise, dimension-$2$ lexicographic, dimension-$2$ balanced mixed-coordinate, and weighted scalar-projection componentwise). Two structural results sit above that stack: a scalar-projection meta-theorem (Theorem 4.13) unifying the projection-based matrix arguments, and a symbolic variable-condition barrier (Theorem 4.14) together with an explicit KO7 KBO corollary (Corollary 4.15) for direct KBO-style comparators. Meta/PumpedBarrierClasses.lean supplies unconditional strengthened-subclass corollaries for the affine, restricted-quadratic, bounded cross-term quadratic, tracked-primary pair, bounded multilinear, max-plus, and projection-based componentwise matrix families. The affine contradiction cutoff is also shown exact for a canonical family in Meta/AffineThresholdSharpness.lean. Meta/BarrierWitness.lean and Meta/BarrierWitness_Extended.lean provide concrete violating triples for the additive and transparent-compositional classes, a packaged affine contradiction once a pump term is supplied, and additional certified extractors for the restricted-quadratic, max-plus, and projected matrix-side pumped subclasses.
@@
C. Certified implementation and artifact engineering. The artifact does more than state theorems. For the guarded SafeStep fragment of KO7 it certifies strong normalization, root-step confluence, a certified normalizer, decidable reachability to safe normal-form targets, partial context-closure strong normalization and confluence, an exact DM-component ordinal calibration at $\omega^{\omega}$, and a triple-lex upper-bound calibration below $\omega^{\omega}\!\cdot\!2$ (Theorems 5.95.11). For the full unguarded Step relation it formalizes two independent root-step termination proofs (nonlinear polynomial and KO7-specialized MPO), a Lean proof of full context-closed strong normalization by lifting the nonlinear witness through contexts, and a concrete derivation-length bound for contextual reductions. The support layer is executable: Meta/BarrierClass_Classifier.lean, Meta/BarrierWitness.lean, and Meta/SynthesisOracle.lean expose a classifier / witness / oracle pipeline for direct-orientation attempts, while Meta/TPDB_Export.lean and lakeexeverifyTpdbExport provide a checked bridge to the external TTT2/CeTA artifact. TTT2 independently reproduces the context-closed termination result; CeTA certifies the external proof.
@@ 3.2 KO7
Table 1: KO7’s 8 kernel rules. R_rec_succ (*) duplicates the step argument $s$.
@@
In full Step, R_eq_refl and R_eq_diff overlap at $\mathrm{eqW}\,a\,a$. The SafeStep fragment adds per-rule guards: $\delta$-flag guards on merge and rec-zero rules serve termination; $\kappa^{M}(a)=\varnothing$ on R_eq_refl and $a\neq b$ on R_eq_diff serve confluence.
@@
SafeStepCtx Partial ctx. Yes (mech.) Yes; exact Newman equivalence n/a
@@
Both targets are full-Step normal forms, so they do not join. The guarded relation SafeStep is introduced to block this overlap and recover unique normal forms for the root ARS.
@@ 4 The orientation barrier
Simple typing by itself is not an escape from the barrier. The artifact includes a simply-typed first-order recursor fragment with separate result, step, and counter sorts, and proves that the additive barrier still survives there. It also proves that the affine barrier survives under an explicit typed step-pump hypothesis, namely, that the step sort still contains arbitrarily large closed terms. The point is structural: typing can block the old untyped pump patterns, but it does not dissolve the obstruction once a typed unbounded step family remains available. Machine-checked in Meta/TypedBarrierSurvival.lean as no_additive_orients_typed_recSucc and no_affine_orients_typed_recSucc_of_stepPump.
Under these conditions, the barrier stack has a clear internal structure rather than a flat catalog. Theorems 4.1 and 4.2 establish the baseline: additive, transparency-constrained compositional, and affine constructor-local measures already fail. Theorems 4.3 to 4.7 then widen the scalar side in stages, asking how much extra nonlinear growth or alternative aggregation can be added before the freeze-and-pump argument breaks; each theorem identifies the surviving obstruction. Theorems 4.9 to 4.13 form the vector / pair block, ending with a single scalar-projection meta-theorem that states the common matrix-side proof shape once. Theorem 4.14 and Corollary 4.15 close the direct symbolic side. This organization matters because the escape results later mirror it: successful methods work only by importing structure outside one of these blocks. Table 2 summarizes the resulting boundary.
@@
Provably blocked classes Twelve direct families are ruled out by theorem: additive and transparent-compositional measures; affine and nonlinear scalar extensions (restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized degree-bounded polynomial, and max-plus); and tracked vector/pair extensions (fixed-dimension componentwise, dimension-$2$ lexicographic, dimension-$2$ balanced mixed-coordinate, and weighted scalar-projection componentwise). The additive and transparent-compositional classes are excluded unconditionally; the remaining ten are excluded under their stated pump or growth hypotheses (Theorems 4.1, 4.2, 4.3, 4.4, 4.5, 4.6, 4.7, 4.9, 4.10, 4.11, 4.12).
@@
Direct out-of-class global orienters The full root relation is oriented internally by the nonlinear polynomial witness and the KO7-specialized MPO, both outside the formalized direct barrier classes (Propositions 4.26, 4.27).
@@ 4.1 Barrier at the schema
Machine-checked in Meta/StepDuplicatingSchema.lean. Computable certificate extractors in Meta/BarrierWitness.lean (additive_witness, compositional_witness, affine_witness) return a concrete violating triple $(b,s,n)$ with a bundled proof of orientation failure: directly from the measure data in the additive and transparent-compositional cases, and from the measure together with a supplied pump term above the required bound in the affine case.
Proof sketch. For Tier 1, set $b=n=\mathrm{base}$ and $s=\mathrm{wrapIter}(w_{\mathrm{succ}})$ (iterated wrapper). By additivity, $M(s)\geq w_{\mathrm{succ}}$. The rewrite right-hand side $\mathrm{wrap}(s,\mathrm{recur}(\mathrm{base},s,\mathrm{base}))$ contributes $w_{\mathrm{wrap}}+M(s)$ beyond the shared recursive subterm, while the left-hand side $\mathrm{recur}(\mathrm{base},s,\mathrm{succ}(\mathrm{base}))$ contributes only $w_{\mathrm{succ}}$. Since $w_{\mathrm{wrap}}\geq 1$ and $M(s)\geq w_{\mathrm{succ}}$, the right-hand side is at least as large, contradicting strict orientation. For Tier 2, set $b=s=n=\mathrm{base}$. By transparency, $c_{\mathrm{succ}}(c_{\mathrm{base}})=c_{\mathrm{base}}$, so $M(\mathrm{recur}(\mathrm{base},\mathrm{base},\mathrm{succ}(\mathrm{base})))=c_{\mathrm{recur}}(c_{\mathrm{base}},c_{\mathrm{base}},c_{\mathrm{base}})$. The right-hand side $c_{\mathrm{wrap}}(c_{\mathrm{base}},c_{\mathrm{recur}}(c_{\mathrm{base}},c_{\mathrm{base}},c_{\mathrm{base}}))$ exceeds the left-hand side by the wrapper-subterm axiom $c_{\mathrm{wrap}}(x,y)>y$. $\square$

### 08 Apr 2026 v8

@@ Abstract
We formalize the orientation boundary for first-order step-duplicating recursors. At the schema level, no measure in twelve formalized direct classes in the barrier package can strictly orient the duplicating step uniformly. These classes include additive and transparent-compositional measures, several scalar growth families, and tracked vector / pair families under stated pump or growth hypotheses. Two meta-results sharpen that stack: a scalar-projection theorem for the matrix-side barriers, and a symbolic variable-condition barrier yielding a KBO-style impossibility corollary for the concrete witness system studied later. To our knowledge, this is the first mechanized object-level barrier theorem, proved without reductions or undecidability arguments, showing that a fixed terminating system lies outside several named direct termination-measure classes. All internal proofs are formalized in Lean 4.
The paper also isolates three escape routes inside the explicit direct universe and extends that analysis on the matrix side by scalar-projection representability. It proves a transparency-essentiality theorem with a concrete witness, a dependency-pair escape characterization, a generalized polynomial boundary theorem isolating failure of frozen base dominance, computable barrier-witness extractors with explicit construction budgets, a coefficient-table classifier, a small synthesis-oracle layer, ablations, and two SCC synchronization families that continue from bounded witness systems to finite cyclic and raw-graph forms. Within this explicit direct universe, every successful orienter succeeds only by importing structural assumptions outside the formalized barrier classes.
The concrete witness calculus KO7 carries a two-layer formal development. For the guarded fragment we internally certify strong normalization, confluence, a certified normalizer, and decidable reachability to safe normal-form targets. We also prove a single-exponential SafeStepCtx derivation-length theorem with a matching explicit lower family, an exact DM-component ordinal calibration at $\omega^{\omega}$, a triple-lex certificate image below $\omega^{\omega}\cdot 2$, and explicit realized lower families in both blocks of that triple-lex image. For the full unguarded system we separately prove root termination by a nonlinear polynomial witness and by a KO7-specialized MPO, lift the nonlinear witness to context-closed strong normalization with an explicit derivation-length bound, and connect the Lean development to TTT2/CeTA through a checked TPDB export, a runnable verifier bridge, and a narrow internal replay of the FAST certificate core.
@@ 1 Introduction
This paper formalizes both sides of this boundary. We prove impossibility theorems at the schema for twelve axiomatized direct measure classes (the barrier package, enumerated below), classify the escape mechanisms, and validate the boundary on a concrete system, KO7, carrying a full certification chain. These twelve classes are not ad hoc: they are the direct scalar and tracked-tuple fragments of the broader tuple-interpretation taxonomy discussed later in Section 4. A separate twelve-family universe drives the escape trichotomy in Section 4; the two are reconciled there.
@@
The observation that additive and KBO-style direct measures cannot orient duplicating rules is classical: KBO cannot orient any duplicating rule because of the variable condition [4], and the dependency pair framework was introduced to handle systems where simplification orderings fail [2]. Middeldorp and Zantema [19] established the hierarchy between simple termination and full termination. Our contribution is to axiomatize a broader transparency-constrained compositional class together with a separate affine/linear constructor-local class, and then prove barrier theorems, an escape characterization, a transparency-essentiality theorem, and a mechanized ablation for those classes. To our knowledge, no prior work formalizes impossibility theorems for named termination-measure classes on a fixed terminating system in the strong object-level sense used here: no member of the specified class can uniformly orient the duplicating step of that system. Mitterwallner, Middeldorp, and Thiemann [20, 28] formalized in Isabelle/HOL the undecidability of orientation by linear polynomial interpretations and by KBO with subterm coefficients. That is a result about the decision problem across all inputs. Their work does not produce standalone impossibility theorems for named systems; ours does. The broader formalization ecosystem (IsaFoR/CeTA, CoLoR) certifies positive termination proofs but not impossibility theorems for measure classes.
The closest neighboring lines of work address different questions. The LICS/AFP undecidability line asks whether a method class is algorithmically decidable across all inputs; certificate ecosystems such as IsaFoR/CeTA and CoLoR ask whether a successful external termination argument can be reconstructed and checked. Our claim is narrower: for one fixed terminating system, we prove object-level impossibility theorems for explicit direct measure families, explain the escape routes, and package constructive witnesses and classifiers around that boundary. The difference is architectural as well as mathematical. The artifact is not another generic certificate checker. It develops a negative boundary theory for one fixed terminating system and then reconnects that Lean development to standard TRS tooling through a checked TPDB export and verifier bridge.
Correspondingly, the paper does not claim that KO7 defeats termination methods in general. The no-go theorems concern the twelve direct measure families of the barrier package and an adjacent direct symbolic barrier. Modular methods and fixed-signature path-order arguments remain on the successful side of the boundary, and the artifact proves exactly that through the DP escape, the nonlinear full-step witness $W$, and the KO7-specialized MPO. KO7 itself is deliberately minimal rather than ad hoc: it is a witness calculus designed to exhibit, in one fixed first-order system, the four structural features that make the boundary stack load-bearing, namely duplication, visible wrappers, a transparent-successor interface on the direct side, and enough auxiliary structure to expose both confluence and certification behavior.
@@
A. The orientation barrier (primary contribution). The main theorem package is a twelve-class direct boundary (the barrier package) for the duplicating step, with the latter ten classes blocked under their stated pump or growth hypotheses and the matrix side unified by a scalar-projection meta-theorem. The count is $2+6+4=12$: two unconditional classes (additive and transparency-constrained compositional), six scalar growth classes (affine, restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized degree-bounded polynomial, and max-plus), and four tracked vector/pair families (fixed-dimension componentwise, dimension-$2$ lexicographic, dimension-$2$ balanced mixed-coordinate, and weighted scalar-projection componentwise). Two structural results sit above that stack: a scalar-projection meta-theorem (Theorem 4.14) unifying the projection-based matrix arguments, and a symbolic variable-condition barrier (Theorem 4.15) together with an explicit KO7 KBO corollary (Corollary 4.16) for direct KBO-style comparators. The generalized bounded-polynomial theorem also now carries a WPO-facing direct polynomial-branch corollary (Corollary 4.7) for orders whose strict comparison is certified by such a bounded polynomial algebra. Meta/PumpedBarrierClasses.lean supplies unconditional strengthened-subclass corollaries for the affine, restricted-quadratic, bounded cross-term quadratic, generalized bounded-polynomial, tracked-primary pair, bounded multilinear, max-plus, weighted-functional projected matrix, and balanced mixed-coordinate sum-pump families. The affine contradiction cutoff is also shown exact for a canonical family in Meta/AffineThresholdSharpness.lean. Meta/BarrierWitness.lean and Meta/BarrierWitness_Extended.lean provide concrete violating triples for the additive and transparent-compositional classes, a packaged affine contradiction once a pump term is supplied, and additional certified extractors for the restricted-quadratic, max-plus, and projected matrix-side pumped subclasses; Meta/BarrierWitness_Budgets.lean packages the corresponding canonical construction budgets, showing that the extracted payload is always base, succIter $k$, or wrapIter $k$ for an explicit computable budget $k$.
@@
C. Certified implementation and artifact engineering. For the guarded SafeStep fragment of KO7, the artifact certifies strong normalization, root-step confluence, a certified normalizer, decidable reachability to safe normal-form targets, partial context-closure strong normalization and confluence, a single-exponential contextual derivation-length theorem with an explicit lower family, an exact DM-component ordinal calibration at $\omega^{\omega}$, and a triple-lex upper-bound calibration below $\omega^{\omega}\!\cdot\!2$ (Theorems 5.95.11). This pins the guarded contextual complexity story at a single-exponential scale up to constant factors in the exponent and isolates the exact $\omega^{\omega}$ order type of the DM component as a standalone calibration result. For the full unguarded Step relation it formalizes two independent root-step termination proofs (nonlinear polynomial and KO7-specialized MPO), a Lean proof of full context-closed strong normalization by lifting the nonlinear witness through contexts, and a concrete derivation-length bound for contextual reductions. The support layer is executable: Meta/BarrierClass_Classifier.lean, Meta/BarrierWitness.lean, and Meta/SynthesisOracle.lean expose a classifier / witness / oracle pipeline for direct-orientation attempts, while Meta/TPDB_Export.lean and lakeexeverifyTpdbExport provide a checked bridge to the external TTT2/CeTA artifact. TTT2 independently reproduces the context-closed termination result; CeTA certifies the external proof.
@@ 3.1 The step-duplicating schema
Meta/TextbookDupInstance.lean packages this case explicitly: the textbook rule $f(x,s(y))\to g(x,f(x,y))$ is turned into a separate schema-level case study, with direct additive, transparent-compositional, and affine barrier corollaries and specialized witness-extractor aliases. The reusable boundary layer therefore applies to a second named duplicating TRS, not only to KO7.
@@ 3.2 KO7
Table 1: KO7’s 8 kernel rules. R_rec_succ (*) duplicates the step argument $s$. In full Step, R_eq_diff has no side condition and applies for all $a,b$.
@@
In full Step, R_eq_refl and R_eq_diff overlap at $\mathrm{eqW}\,a\,a$. The SafeStep fragment adds per-rule guards: $\delta$-flag guards on the merge-void and rec-zero rules serve termination; the side conditions $\delta\text{-flag}(t)=0$ and $\kappa^{M}(t)=\varnothing$ on R_merge_cancel control the collapsing merge branch; and $\kappa^{M}(a)=\varnothing$ on R_eq_refl together with $a\neq b$ on R_eq_diff serve confluence.
@@
Table 2: Certified status of the primary KO7 reduction relations.
@@
SafeStepCtx Partial ctx. Yes (mech.) Yes; Newman characterization retained n/a
@@
Both targets are full-Step root normal forms, so they do not join. The guarded relation SafeStep is introduced to block this overlap and recover unique normal forms for the root ARS.
@@ 4 The orientation barrier
Section preview.
This section isolates a concrete frontier between direct measure classes that provably fail and methods that succeed by importing structure outside those classes. Inside the formalized direct universe, the duplicating step is excluded by the schema, affine, restricted-quadratic, bounded cross-term, bounded multilinear, generalized bounded-polynomial, max-plus, tracked componentwise / lexicographic / balanced mixed-coordinate / weighted scalar-projection vector and pair barriers, together with the symbolic variable-condition (KBO) corollary. Crossing the frontier requires one of three moves exhibited in the artifact: projection-style transformation (dependency pairs), direct nonlinear escape outside the formalized classes (the global witness $W$ and the KO7-specialized MPO), or ablation of the offending structure itself. A more detailed recap appears at the end of the section as the Frontier summary.
@@
Simple typing by itself is not an escape from the barrier. The artifact includes a simply-typed first-order recursor fragment with separate result, step, and counter sorts, and proves that the additive barrier still survives there. It also proves that the affine barrier survives under an explicit typed step-pump hypothesis, namely, that the step sort still contains arbitrarily large closed terms. The point is structural: typing can block the old untyped pump patterns, but it does not dissolve the obstruction once a typed unbounded step family remains available. The Lean proofs are no_additive_orients_typed_recSucc and no_affine_orients_typed_recSucc_of_stepPump in Meta/TypedBarrierSurvival.lean. The same sort-indexed syntax is also repackaged in Meta/ManySortedBarrierSurvival.lean as a specialized many-sorted first-order presentation, yielding no_additive_orients_manySorted_recSucc and no_affine_orients_manySorted_recSucc_of_stepPump. This is a theorem-backed many-sorted extension of the KO7 recursor fragment, not a full mechanization of the general Aoto–Yamada translation theorem [1].
Under these conditions, the barrier stack has a clear internal structure rather than a flat catalog. Theorems 4.1 and 4.2 establish the baseline: additive, transparency-constrained compositional, and affine constructor-local measures already fail. Theorems 4.3 to 4.8 then widen the scalar side in stages, asking how much extra nonlinear growth or alternative aggregation can be added before the freeze-and-pump argument breaks; each theorem identifies the surviving obstruction. Theorems 4.10 to 4.14 form the vector / pair block, ending with a single scalar-projection meta-theorem that states the common matrix-side proof shape once. Theorem 4.15 and Corollary 4.16 close the direct symbolic side. This organization matters because the escape results later mirror it: successful methods work only by importing structure outside one of these blocks. Table 3 summarizes the resulting boundary.
@@
Provably blocked classes Twelve direct families are ruled out by theorem: additive and transparent-compositional measures; affine and nonlinear scalar extensions (restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized degree-bounded polynomial, and max-plus); and tracked vector/pair extensions (fixed-dimension componentwise, dimension-$2$ lexicographic, dimension-$2$ balanced mixed-coordinate, and weighted scalar-projection componentwise). The additive and transparent-compositional classes are excluded unconditionally; the remaining ten are excluded under their stated pump or growth hypotheses (Theorems 4.1, 4.2, 4.3, 4.4, 4.5, 4.6, 4.8, 4.10, 4.11, 4.12, 4.13).
@@
Direct out-of-class global orienters The full root relation is oriented internally by the nonlinear polynomial witness and the KO7-specialized MPO, both outside the formalized direct barrier classes (Propositions 4.27, 4.28).
Table 4: Status of the main barrier side conditions.
Class block Status Explanation
Additive; transparent-compositional Unconditional No extra growth or dominance side condition is needed.
Affine; restricted quadratic; max-plus; tracked-primary vector / pair; weighted scalar-projection; balanced mixed-coordinate Standard pump hypotheses The required growth hypotheses are easy to discharge from a successor pump or wrap/base pump and are internalized as strengthened subclasses where appropriate in Meta/PumpedBarrierClasses.lean.
Bounded cross-term quadratic; bounded multilinear; generalized bounded polynomial Load-bearing dominance hypotheses These classes carry genuine frozen-coefficient or base-dominance constraints that delimit the theorem scope rather than following automatically from the schema alone.
This summary is only a status map for the stated hypotheses, not an extra theorem. Its purpose is to separate routine pump side conditions from the genuinely restrictive dominance assumptions that remain part of the mathematical boundary.
@@ 4.1 Barrier at the schema
The Lean proof is in Meta/StepDuplicatingSchema.lean. Computable certificate extractors in Meta/BarrierWitness.lean (additive_witness, compositional_witness, affine_witness) return a concrete violating triple $(b,s,n)$ with a bundled proof of orientation failure: directly from the measure data in the additive and transparent-compositional cases, and from the measure together with a supplied pump term above the required bound in the affine case. The companion module Meta/BarrierWitness_Budgets.lean packages the corresponding construction-budget statements: for the current extractor layer, the generated step payload is always the base term, a successor chain succIterk, or a wrapper chain wrapIterk, with an explicit budget $k$ read off from the coefficient data.
Argument. For Tier 1, set $b=n=\mathrm{base}$ and $s=\mathrm{wrapIter}(w_{\mathrm{succ}})$ (iterated wrapper). By additivity, $M(s)\geq w_{\mathrm{succ}}$. The rewrite right-hand side $\mathrm{wrap}(s,\mathrm{recur}(\mathrm{base},s,\mathrm{base}))$ contributes $w_{\mathrm{wrap}}+M(s)$ beyond the shared recursive subterm, while the left-hand side $\mathrm{recur}(\mathrm{base},s,\mathrm{succ}(\mathrm{base}))$ contributes only $w_{\mathrm{succ}}$. Since $w_{\mathrm{wrap}}\geq 1$ and $M(s)\geq w_{\mathrm{succ}}$, the right-hand side is at least as large, contradicting strict orientation. For Tier 2, set $b=s=n=\mathrm{base}$. By transparency, $c_{\mathrm{succ}}(c_{\mathrm{base}})=c_{\mathrm{base}}$, so $M(\mathrm{recur}(\mathrm{base},\mathrm{base},\mathrm{succ}(\mathrm{base})))=c_{\mathrm{recur}}(c_{\mathrm{base}},c_{\mathrm{base}},c_{\mathrm{base}})$. The right-hand side $c_{\mathrm{wrap}}(c_{\mathrm{base}},c_{\mathrm{recur}}(c_{\mathrm{base}},c_{\mathrm{base}},c_{\mathrm{base}}))$ exceeds the left-hand side by the wrapper-subterm axiom $c_{\mathrm{wrap}}(x,y)>y$. We retain the symmetric pair of wrapper-subterm axioms in the theorem interface for standardity and for downstream uniformity across the compositional API, although only the second inequality drives the contradiction. $\square$